OpenStudy (dls):

The points on the curve 9y^2=x^3 where the normal to the curve makes equal intercepts with the axes=?

OpenStudy (dls):

@hartnn

hartnn (hartnn):

did u differentiate to get dy/dx ?

OpenStudy (dls):

yes

OpenStudy (dls):

\[\LARGE 18yy'=3x^2=>y'=\frac{3x^2}{18y}=\frac{x^2}{6y}\]

hartnn (hartnn):

the slope of the normal will be \(\Large \dfrac{-1}{\dfrac{dy}{dx}}\)

OpenStudy (dls):

normal slope.. \[\LARGE \frac{-6y}{x^2}\]

OpenStudy (dls):

not getting how to "apply" the 2nd part of the question

hartnn (hartnn):

let the point be x1, y1 make an equation of line since you have point and slope

hartnn (hartnn):

then find out the intercepts, equate them

hartnn (hartnn):

you know how to find intercepts, right ?

OpenStudy (dls):

\[\large y-y_1 =\frac{-6y}{x^2} (x-x_1)\] umm put x=0 to get y intercept and similarly for x?

hartnn (hartnn):

exactly! :)

OpenStudy (dls):

then ill get (0,0) ?

OpenStudy (dls):

both intercepts 0? :/ how is that possible

hartnn (hartnn):

ok, wait, i think what we need to do is, express that line equation in the form of x/a +y/b=1 then a and b are intercepts

OpenStudy (dls):

this is a curve but :O

OpenStudy (dls):

oh okay

hartnn (hartnn):

equating a and b , you will get one equation. the other equation is the fact that x1,y1 lies on the curve 9y1^2 = x1^3

OpenStudy (dls):

hogaya :D thanks!

hartnn (hartnn):

welcome ^_^